# PLRD: Partially Linear Regression Discontinuity Inference

### v1

@@ A.4 Ablation analysis
    Bias-aware intervals for local linear regression following Armstrong and Kolesár (2020):
@@
      Build the confidence interval $\widehat{\tau}_{\texttt{LLR}}\pm 2.181\cdot\widehat{\text{SE}}_{\texttt{LLR}}$ (Armstrong and Kolesár, 2020).
@@
    Bias-aware intervals for local quadratic regression following Armstrong and Kolesár (2020):
@@
      Fit a local quadratic regression using the rdrobust package (i.e., with $p=2$ in Section A.1), with default bandwidth selection rule.
@@
      Obtain the LQR estimator $\widehat{\tau}_{\texttt{LQR}}$ (defined as $\widehat{\tau}_{2}(h_{n})$ in Section A.1) with its conventional standard error estimate $\widehat{\text{SE}}_{\texttt{LQR}}$
@@
      Build the confidence interval $\widehat{\tau}_{\texttt{LQR}}\pm 2.113\cdot\widehat{\text{SE}}_{\texttt{LQR}}$ (Armstrong and Kolesár, 2020).
@@
    Algorithmically, this is similar to our Algorithm 1; we have also implemented this as a part of our plrd package.
@@
    Algorithmically, this is similar to the optrdd package by Imbens and Wager (2019). We also provide an implementation of this method as a part of our replication files.
Dataset Alt. (a) Alt. (b) Alt. (c) Alt. (d) plrd
Lee (2008) 0.040 0.047 0.041 0.038 0.035
(96.5%) (96.3%) (97.8%) (99.8%) (98.3%)
Cattaneo et al. (2015) 5.956 7.267 5.259 9.376 4.525
(96.6%) (96.4%) (97.3%) (99.7%) (97.3%)
Matsudaira (2008) (read.) 0.068 0.070 0.071 0.084 0.060
(96.1%) (96.7%) (97.1%) (96.8%) (97.1%)
Matsudaira (2008) (math) 0.050 0.056 0.059 0.073 0.051
(94.4%) (95.7%) (97.3%) (95.2%) (97.0%)
Jacob and Lefgren (2004) (read.) 0.075 0.092 0.078 0.075 0.065
(96.8%) (96.5%) (96.0%) (99.3%) (94.6%)
Jacob and Lefgren (2004) (math) 0.087 0.102 0.082 0.077 0.069
(95.6%) (94.9%) (97.2%) (96.2%) (95.0%)
Oreopoulos (2006) 0.218 0.401 0.225 0.185 0.194
(95.2%) (96.9%) (98.6%) (94.5%) (96.1%)
Ludwig and Miller (2007) 4.516 5.564 4.075 3.846 3.542
(97.2%) (96.5%) (97.6%) (98.2%) (97.4%)
Meyersson (2014) 7.079 8.796 5.905 7.837 5.095
(95.7%) (94.7%) (96.7%) (98.5%) (97.3%)
Pure noise simulation 1.098 1.375 1.612 1.179 0.847
(96.6%) (98.6%) (97.3%) (98.2%) (96.3%)
Simulation Setting 1 0.223 0.278 0.276 0.202 0.235
(92.7%) (97.5%) (97.2%) (92.5%) (94.6%)
Simulation Setting 2 0.297 0.289 0.375 0.493 0.309
(86.4%) (86.1%) (97.7%) (98.2%) (51.5%)
Simulation Setting 3 0.236 0.279 0.280 0.300 0.239
(93.3%) (96.2%) (97.1%) (98.2%) (95.9%)
Simulation Setting 4 0.223 0.276 0.279 0.255 0.236
(93.6%) (96.5%) (96.7%) (98.1%) (97.5%)
Table 5: Comparison of PLRD with other alternatives: first two columns correspond to inflating conventional confidence intervals: instead of $1.96$, we multiply the standard error by a factor of (a) $2.181$ or (b) $2.113$, as suggested by Armstrong and Kolesár (2020); the next two columns correspond to (c) PLRD minus Assumption 2, and (d) PLRD minus Assumption 1; the last column corresponds to our PLRD method. We write the empirical coverages in parentheses and the average width in the rows. Other details of this table are same as for Tables 1, 2 and 4 in the main text.
The empirical results demonstrate that the rule-of-thumb methods (AK20 (a) and AK20 (b)) achieve valid coverage but at the cost of wider intervals, mirroring our earlier findings in Tables 1, 2 and 4. The relaxed variants of PLRD offer insights into the role of each assumption: Dropping either Assumption 1 or Assumption 2 generally leads to more conservative intervals, suggesting that both assumptions contribute to the efficiency of PLRD.
Appendix B Additional details on our real data experiments
Here we provide detailed descriptions of the real datasets analyzed in the main paper. We reference each dataset by the paper that introduced it, same as in Tables 1, 3 and 5.
@@ Appendix B Additional details on our real data experiments
    Lee (2008): This dataset, frequently used for illustrating regression discontinuity methods, examines the incumbency advantage in U.S. House elections. The running variable is the margin of victory, defined as the difference in vote share between the Democratic candidate and their strongest opponent, with the threshold set at $c=0$. The outcome of interest is the probability of winning subsequent elections. The dataset consists of $6,558$ observations from U.S. House elections between 1946 and 1998. Lee (2008) reported a significant positive effect of incumbency on the probability of winning subsequent elections, indicating an incumbency advantage; see Table 1 and Figure 2 of Lee (2008) for more details.
@@
    Cattaneo et al. (2015): This is similar to Lee (2008) except that they study the incumbency advantage in U.S. Senate elections instead of House elections. The running variable is the margin of victory, defined as the difference in vote percentages (ranging from $-100$ to $100$) between the Democratic candidate and their strongest opponent, with the threshold set at $c=0$. The outcome of interest is the percentage of winning subsequent elections. The dataset consists of $1,390$ observations ($1,297$ excluding missing outcomes) from U.S. House elections between 1914 and 2010. Our findings align with the original paper Cattaneo et al. (2015) as well as the replication in Cattaneo et al. (2024).
@@
    Matsudaira (2008): This dataset records fifth-grade students’ math and reading scores in an unspecified large school district in the northeast from 2001 to 2002. Matsudaira (2008) examined the impact of mandatory summer school on students’ performance. The running variable is the minimum of students’ test scores in math and reading, with a score below the passing threshold for either math or reading determining mandatory attendance in the summer school. The primary outcomes are post-summer school reading and math test scores. We use a version of this dataset shared by the author with K. Kalyanamaran and G. Imbens in 2008. This dataset has a slightly larger sample size $n=68,798$ than the one used in the original paper ($n=66,839$). Our findings mirror the results in Matsudaira (2008, Table 4) for fifth-grade students.
@@
    Jacob and Lefgren (2004): This dataset contains the record of third-grade student achievement in Chicago Public Schools from 1997 to 1999. Jacob and Lefgren (2004) investigated the effectiveness of remedial education programs on student achievement, using a regression discontinuity design based on test score. We use the same dataset as Imbens and Zajonc (2009), which contains $70,831$ observations after dropping entries with missing information. The running variable is the minimum of the math and reading score for an exam in the spring and the outcome is the post-remedial program reading (resp. math) test scores, reported as Rasch test scores rather than grade equivalents. Rasch test scores are normalized scores (zero mean, unit s.d.) for the population of third-grade students from 1997 to 1999. Our results are aligned with the findings in Jacob and Lefgren (2004) for third-grade students.
@@
    Oreopoulos (2006): This study concerns the impact of compulsory schooling laws on educational attainment. The running variable is the individual’s year of birth, with thresholds determined by changes in compulsory schooling laws. The outcome here is years of completed education. Our dataset comprises $73,954$ observations, differing slightly from the original study due to variations in preprocessing; again, we use the same dataset also used in Kolesár and Rothe (2018) and focus on the reduced form analysis rather than a fuzzy RD approach (see Table 1, Panel C of the corrigendum to Oreopoulos (2006)). Our findings are in line with Table 2 of Kolesár and Rothe (2018).
@@
    Ludwig and Miller (2007): This study examines the effectiveness of the Head Start program on children’s health and schooling. The running variable is the 1960 poverty rate, and the threshold corresponds to the qualifying cutoff for program participation. Following Calonico et al. (2014), we consider as outcome variable the mortality rates per 100,000 for children between 5 and 9 years old, with Head Start-related causes, for 1973–1983. After cleaning missing values, this dataset includes $2,783$ observations. Table III and Figure IV (Panel A) in Ludwig and Miller (2007) illustrate that the Head Start program had a negative impact on the outcomes.
@@
    Meyersson (2014): This study investigates the effect of Islamic party control in local governments on women’s rights, in particular on the educational attainment of young women, focusing on Turkey’s 1994 mayoral elections. The study employs an RD design where the running variable is the Islamic margin of victory—the vote share gap between the leading Islamic and secular parties, thus with a natural cutoff of zero. The analysis is conducted at the municipality level and has $2,629$ observations. We replicate part of the study where the primary outcome is the school attainment for women who were (potentially) in high school during the period 1994-2000, measured with variables extracted from the 2000 census. This data have been used as a running example in Cattaneo et al. (2024). Our results align with the illustrations in Cattaneo et al. (2024) and the original paper Meyersson (2014) — in municipalities where the Islamic party barely won, the educational attainment of women is roughly 3% higher than in municipalities where the party barely lost.
Appendix C Implementation via Convex Optimization
@@
$\mathbb{E}\left[\sum_{i=1}^{n}\gamma_{i}Y_{i}\mid\{X_{i}\}_{i=1}^{n}\right]-\tau_{c}$ $=\sum_{i=1}^{n}\gamma_{i}\mu_{W_{i}}(X_{i})-\tau_{c}$
@@
[…]
@@
$(\widehat{\gamma},\widehat{t}\,):=$ $\mathop{\rm{minimize}}_{\gamma,t}\left\{B^{2}t^{2}+\sum_{i=1}^{n}\gamma_{i}^{2}\sigma_{i}^{2}\right\}$
@@
The function $\widetilde{\rho}(\cdot)$ in the above formulation corresponds to the function $\lambda_{1}\rho(\cdot)$ in the original problem. Thus, we can recover the optimal variables $(\widehat{\gamma},\widehat{t}\,)$ for the primal problem (24) from the dual optimal variable $\widehat{\theta}$ (i.e., solution of (25)), as follows.
@@
Appendix D Main proofs
@@
For simplicity in notation, we assume without loss of generality that the regression discontinuity threshold is $c=0$. We split the proof into some auxiliary results, namely, Lemmas 10, 8, 7, 9 and 11 in Appendix E. First, using tools from empirical process theory (see van der Vaart and Wellner (1996) for a book-level treatment), we prove a uniform concentration bound for the loss functions of the empirical and population dual problems. Recall that the dual parameter space is defined as
@@
Appendix E Auxiliary results
[…]
@@ Appendix E Auxiliary results
[…]
@@
In Algorithm 1, we sample splitting to estimate the unknown Lipschitz constant $B$ and the homoscedastic variance $\sigma^{2}={\rm Var}(Y_{i}\mid X_{i})$. However, for ease of exposition, we assume throughout this proof that an estimate $\widehat{B}$ of $B$ and an estimate $\widehat{\sigma}^{2}$ of $\sigma^{2}$ is available independently of the data using which we solve the empirical dual problem. We define
@@
[…]
@@
[…]
@@
[…]
@@
Equipped with Eqs. 34, 35 and 36, we are ready to simplify $L_{n}(\theta)-L_{n}(\theta_{n}^{*})=\widetilde{L}_{n}(\widetilde{\rho},\lambda)-\widetilde{L}_{n}(\widetilde{\rho}_{n}^{*},\lambda_{n}^{*})$, as follows.
@@
Consider the neighborhood $\Theta_{n}(\ell,\delta,\varepsilon)$ of the optimizer $\theta_{n}^{*}=\theta_{n}^{*}(\widehat{\sigma}^{2},B)$ of the population dual problem, as defined in Lemma 7. Denote by $\widehat{\theta}_{n}=\widehat{\theta}_{n}(\widehat{\sigma}^{2},\widehat{B})$ the optimizer of the empirical dual problem. Assume that $\widehat{B}\rightarrow_{p}B$, and that $\ell=o(n^{-1/12})$. For any $\delta,\eta\in(0,1)$, and $\varepsilon=O(n^{-1/4})$, it holds that
@@
for all sufficiently large $n$. Finally, note that since the empirical loss function $\widehat{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})$ is convex in $\theta$, the above event implies that $\widehat{\theta}_{n}=\mathop{\rm argmin}_{\theta}\widehat{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})$ must be inside $\widetilde{\Theta}_{n}$. Thus,
@@
[…]
@@
$(\widehat{\gamma}_{n},\widehat{t}_{n}):=\mathop{\rm argmin}_{\gamma,\,t}\left(\widehat{B}^{2}t^{2}+\widehat{\sigma}^{2}\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\right)\quad\text{subject to}\quad\sup_{\rho\in\mathcal{R}}\sum_{i=1}^{n}\gamma_{i}\rho(X_{i})\leq t,$
@@
[…]
@@
[…]

### v2

@@ B.4 Ablation Analysis
Dataset Alt. (a) Alt. (b) Alt. (c) Alt. (d) plrd
Lee (2008) 0.041 0.051 0.045 0.043 0.045
(96.6%) (95.8%) (98.2%) (99.7%) (98.1%)
Cattaneo et al. (2015) 5.567 6.880 5.652 9.709 3.870
(96.6%) (96.0%) (98.3%) (99.8%) (96.1%)
Ludwig and Miller (2007) 6.626 8.104 7.049 3.905 4.272
(96.8%) (96.3%) (97.8%) (87.4%) (96.8%)
Meyersson (2014) 7.011 8.736 6.432 7.599 4.618
(96.3%) (95.9%) (96.7%) (97.9%) (96.3%)
Matsudaira (2008) (read.) 0.056 0.070 0.059 0.085 0.059
(90.5%) (95.4%) (97.7%) (94.8%) (97.7%)
Matsudaira (2008) (math) 0.051 0.061 0.052 0.075 0.052
(95.0%) (95.1%) (97.6%) (92.4%) (97.6%)
Jacob and Lefgren (2004) (read.) 0.076 0.099 0.082 0.070 0.076
(96.4%) (95.1%) (95.4%) (97.8%) (97.5%)
Jacob and Lefgren (2004) (math) 0.066 0.085 0.070 0.070 0.061
(96.9%) (95.0%) (95.5%) (97.1%) (97.9%)
Oreopoulos (2006) 0.210 0.393 0.244 0.183 0.124
(96.2%) (96.6%) (98.3%) (98.3%) (96.2%)
Pure noise simulation 1.611 2.053 1.678 1.049 1.011
(96.2%) (98.4%) (97.1%) (98.3%) (97.2%)
Simulation Setting 1 0.238 0.301 0.250 0.169 0.237
(93.5%) (97.5%) (95.2%) (88.9%) (94.8%)
Simulation Setting 2 0.330 0.313 0.389 0.468 0.389
(87.9%) (86.6%) (97.3%) (83.9%) (97.3%)
Simulation Setting 3 0.252 0.300 0.290 0.421 0.290
(94.9%) (97.5%) (99.6%) (99.9%) (99.6%)
Simulation Setting 4 0.235 0.300 0.276 0.368 0.178
(96.2%) (98.2%) (99.3%) (100.0%) (96.1%)
Table 4: Comparison of PLRD with other alternatives: first two columns correspond to inflating conventional confidence intervals: instead of $1.96$, we multiply the standard error by a factor of (a) $2.181$ or (b) $2.113$, as suggested by Armstrong and Kolesár (2020); the next two columns correspond to (c) PLRD minus Assumption 2, and (d) PLRD minus Assumption 1; the last column corresponds to our PLRD method. We write the empirical coverages in parentheses and the average width in the rows. Other details of this table are same as for Tables 1 and 3 in the main text.
@@
    Bias-aware intervals for local linear regression with inflated critical values as proposed by Armstrong and Kolesár (2020): [^17] Our end-to-end implementation differs from that in Armstrong and Kolesár (2020), as we choose bandwidths for local linear regression using the bandwidth selector of Imbens and Kalyanaraman (2012) as implemented in rdrobust; whereas Armstrong and Kolesár (2020) discuss adaptive bandwidth choice using an alternative rule of thumb. Here, we chose to pick bandwidths as in rdrobust to enable a cleaner comparison to the conventional estimator.
@@
      Build the confidence interval $\widehat{\tau}_{\texttt{LLR}}\pm 2.181\cdot\widehat{\text{SE}}_{\texttt{LLR}}$.
@@
    Bias-aware intervals for local quadratic regression with inflated critical values as proposed by Armstrong and Kolesár (2020):
@@
      Fit a local quadratic regression using the rdrobust package (i.e., with $p=2$ in Section B.1), with default bandwidth selection rule.
@@
      Obtain the LQR estimator $\widehat{\tau}_{\texttt{LQR}}$ (defined as $\widehat{\tau}_{2}(h_{n})$ in Section B.1) with its conventional standard error estimate $\widehat{\text{SE}}_{\texttt{LQR}}$
@@
      Build the confidence interval $\widehat{\tau}_{\texttt{LQR}}\pm 2.113\cdot\widehat{\text{SE}}_{\texttt{LQR}}$.
@@
    Algorithmically, this is similar to our Algorithm 1; we also implemented this as a part of our plrd package.
@@
    Algorithmically, this is similar to the optrdd package by Imbens and Wager (2019). [^18] In fact, over the weaker smoothness class $\mathcal{M}_{\text{H2},\,M}$, Assumption 2 on partial linearity becomes much less useful, because the second derivative of the baseline effect can have (bounded) sudden jumps anyways. As such, we also expect numerical performance of this ablation baseline to be close to optrdd. We also provide an implementation of this method as a part of our replication files.
The empirical results demonstrate that the rule-of-thumb methods (AK20 (a) and AK20 (b)) achieve valid coverage but at the cost of wider intervals, mirroring our earlier findings in Tables 1 and 3. The relaxed variants of PLRD offer insights into the role of each assumption: Dropping either Assumption 1 or Assumption 2 generally leads to more conservative intervals, suggesting that both assumptions contribute to the efficiency of PLRD.
Appendix C Additional Details on Real Data Experiments
Here we provide detailed descriptions of the real datasets analyzed in the main paper. We reference each dataset by the paper that introduced it, same as in Tables 1, 2 and 4.
@@ Appendix C Additional Details on Real Data Experiments
    Lee (2008): This dataset, frequently used for illustrating regression discontinuity methods, examines the incumbency advantage in U.S. House elections. The running variable is the margin of victory, defined as the difference in vote share between the Democratic candidate and their strongest opponent, with the threshold set at $c=0$. The outcome of interest is the probability of winning subsequent elections. The dataset consists of 6,5586{,}558 observations from U.S. House elections between 1946 and 1998. Lee (2008) reported a significant positive effect of incumbency on the probability of winning subsequent elections, indicating an incumbency advantage; see Table 1 and Figure 2 of Lee (2008) for more details.
@@
    Cattaneo et al. (2015): This is similar to Lee (2008) except that they study the incumbency advantage in U.S. Senate elections instead of House elections. The running variable is the margin of victory, defined as the difference in vote percentages (ranging from 100-100 to $100$) between the Democratic candidate and their strongest opponent, with the threshold set at $c=0$. The outcome of interest is the percentage of winning subsequent elections. The dataset consists of 1,3901{,}390 observations (1,2971{,}297 excluding missing outcomes) from U.S. House elections between 1914 and 2010. Our findings align with the original paper Cattaneo et al. (2015) as well as the replication in Cattaneo et al. (2024).
@@
    Matsudaira (2008): This dataset records fifth-grade students’ math and reading scores in an unspecified large school district in the northeast from 2001 to 2002. Matsudaira (2008) examined the impact of mandatory summer school on students’ performance. The running variable is the minimum of students’ test scores in math and reading, with a score below the passing threshold for either math or reading determining mandatory attendance in the summer school. The primary outcomes are post-summer school reading and math test scores. We use a version of this dataset shared by the author with K. Kalyanamaran and G. Imbens in 2008. This dataset has a slightly larger sample size n=68,798n=68{,}798 than the one used in the original paper (n=66,839n=66{,}839). Our findings mirror the results in Matsudaira (2008, Table 4) for fifth-grade students.
@@
    Jacob and Lefgren (2004): This dataset contains the record of third-grade student achievement in Chicago Public Schools from 1997 to 1999. Jacob and Lefgren (2004) investigated the effectiveness of remedial education programs on student achievement, using a regression discontinuity design based on test score. We use the same dataset as Imbens and Zajonc (2009), which contains 70,83170{,}831 observations after dropping entries with missing information. The running variable is the minimum of the math and reading score for an exam in the spring and the outcome is the post-remedial program reading (resp. math) test scores, reported as Rasch test scores rather than grade equivalents. Rasch test scores are normalized scores (zero mean, unit s.d.) for the population of third-grade students from 1997 to 1999. Our results are aligned with the findings in Jacob and Lefgren (2004) for third-grade students.
@@
    Oreopoulos (2006): This study concerns the impact of compulsory schooling laws on educational attainment. The running variable is the individual’s year of birth, with thresholds determined by changes in compulsory schooling laws. The outcome here is years of completed education. Our dataset comprises 73,95473{,}954 observations, differing slightly from the original study due to variations in preprocessing; again, we use the same dataset also used in Kolesár and Rothe (2018) and focus on the reduced form analysis rather than a fuzzy RD approach (see Table 1, Panel C of the corrigendum to Oreopoulos (2006)). Our findings are in line with Table 2 of Kolesár and Rothe (2018).
@@
    Ludwig and Miller (2007): This study examines the effectiveness of the Head Start program on children’s health and schooling. The running variable is the 1960 poverty rate, and the threshold corresponds to the qualifying cutoff for program participation. Following Calonico et al. (2014), we consider as outcome variable the mortality rates per 100,000 for children between 5 and 9 years old, with Head Start-related causes, for 1973–1983. After cleaning missing values, this dataset includes 2,7832{,}783 observations. Table III and Figure IV (Panel A) in Ludwig and Miller (2007) illustrate that the Head Start program had a negative impact on the outcomes.
@@
    Meyersson (2014): This study investigates the effect of Islamic party control in local governments on women’s rights, in particular on the educational attainment of young women, focusing on Turkey’s 1994 mayoral elections. The study employs an RD design where the running variable is the Islamic margin of victory—the vote share gap between the leading Islamic and secular parties, thus with a natural cutoff of zero. The analysis is conducted at the municipality level and has 2,6292{,}629 observations. We replicate part of the study where the primary outcome is the school attainment for women who were (potentially) in high school during the period 1994–2000, measured with variables extracted from the 2000 census. This data have been used as a running example in Cattaneo et al. (2024). Our results align with the illustrations in Cattaneo et al. (2024) and the original paper Meyersson (2014) — in municipalities where the Islamic party barely won, the educational attainment of women is roughly 3% higher than in municipalities where the party barely lost.
Appendix D Implementation via Convex Optimization
@@
$\mathbb{E}\left[\sum_{i=1}^{n}\gamma_{i}Y_{i}\,\bigg|\,\{X_{i}\}_{i=1}^{n}\right]-\tau_{c}$ $=\sum_{i=1}^{n}\gamma_{i}\mu_{W_{i}}(X_{i})-\tau_{c}$
@@
Note that for any $\mu_{w}(\cdot)\in\mathcal{M}_{B}$, the function $\rho(\cdot)$ belongs to the function class $\mathcal{R}:=\{f:f(0)=f^{\prime}(0)=f^{\prime\prime}(0)=0,f^{\prime\prime}\text{ is }1\text{-Lipschitz}\}$. We can therefore rewrite the optimization problem (25) as:
@@
[…]
@@
[…]
@@
Appendix E Main Proofs
@@
For simplicity in notation, we assume without loss of generality that the regression discontinuity threshold is $c=0$. We split the proof into some auxiliary results, namely, Lemmas 11, 9, 8, 10 and 12 in Appendix F. First, using tools from empirical process theory (see van der Vaart and Wellner (1996) for a book-level treatment), we prove a uniform concentration bound for the loss functions of the empirical and population dual problems. Recall that the dual parameter space is defined as
@@
Appendix F Auxiliary Results
To set the stage, we recall the key concepts of entropy with and without bracketing for function classes, which will be central to the results presented in this section. Consider a class of real-valued functions $\mathcal{F}$ defined on a set $\mathcal{Z}$, equipped with a norm $\|\cdot\|$. The covering number $N(\varepsilon,\mathcal{F},\|\cdot\|)$ is the minimum number of balls of radius $\varepsilon$ needed to cover $\mathcal{F}$. The centers of the covering balls need not belong to $\mathcal{F}$, but they are required to have finite norms. The entropy (without bracketing) is defined as the logarithm of the covering number. An $\varepsilon$-bracket is a pair of functions $(l,u)$ such that $\|u-l\|<\varepsilon$ and $l\leq f\leq u$ for all $f\in\mathcal{F}$. The bracketing number $N_{[\,]}(\varepsilon,\mathcal{F},\|\cdot\|)$ is the minimum number of $\varepsilon$-brackets needed to cover $\mathcal{F}$. The upper and lower bounds $u$ and $l$ of the brackets need not belong to $\mathcal{F}$, but they are required to have finite norms. The entropy with bracketing is defined as the logarithm of the bracketing number.
@@ Appendix F Auxiliary Results
$\mathcal{F}_{B,\,\ell,\,k}=\{\rho\mid\rho:[-\ell,\ell]\to\mathbb{R},\ \rho(0)=\rho^{\prime}(0)=\cdots=\rho^{(k)}(0)=0,\ \rho^{(k)}\text{ is $B$-Lipschitz}\},$
@@
In Algorithm 1, we sample splitting to estimate the unknown Lipschitz constant $B$ and the homoskedastic variance $\sigma^{2}={\rm Var}(Y_{i}\mid X_{i})$. However, for ease of exposition, we assume throughout this proof that an estimate $\widehat{B}$ of $B$ and an estimate $\widehat{\sigma}^{2}$ of $\sigma^{2}$ is available independently of the data using which we solve the empirical dual problem. We define
@@
$\lesssim\frac{1}{\sqrt{n}}\int_{0}^{\|F\|_{P,2}}\sqrt{\log N_{[\,]}(\varepsilon^{\prime},\mathcal{A},L_{2}(P))}\,d\varepsilon^{\prime}$
$\lesssim\frac{1}{\sqrt{n}}\int_{0}^{F}(\varepsilon^{\prime})^{-1/4}\left(\lambda_{n,1}^{*}\ell^{3}\right)^{1/2}\,d\varepsilon^{\prime}$
@@
$\mathbb{E}\sup_{(\rho,\lambda)\in\Theta_{n}}|(\mathbb{P}_{n}-\mathbb{P})(\lambda_{-1}^{\top}v(X,W))^{2}|$ $\lesssim\frac{1}{\sqrt{n}}\int_{0}^{16\|\lambda_{n,-1}^{*}\|^{2}}\sqrt{\log N_{[\,]}(\varepsilon^{\prime},\Lambda_{n},L_{2}(P))}\,d\varepsilon^{\prime}$
@@
$\mathbb{E}\sup_{(\rho,\lambda)\in\Theta_{n}}|(\mathbb{P}_{n}-\mathbb{P})\rho(X)\lambda_{-1}^{\top}v(X,W)|$ $\lesssim\frac{1}{\sqrt{n}}\int_{0}^{F_{ab}}\left((\varepsilon^{\prime})^{-1/6}F_{ab}^{1/6}+(\varepsilon^{\prime})^{-1/2}\varepsilon^{1/2}F_{ab}^{1/2}\right)\,d\varepsilon^{\prime}$
@@
Equipped with Equations 36, 37 and 38, we are ready to simplify $L_{n}(\theta)-L_{n}(\theta_{n}^{*})=\widetilde{L}_{n}(\widetilde{\rho},\lambda)-\widetilde{L}_{n}(\widetilde{\rho}_{n}^{*},\lambda_{n}^{*})$, as follows.
@@
[…]
@@
[…]
@@
Recall that $\widehat{\lambda}_{n,1}(\widehat{\sigma}^{2},\widehat{B})=2\widehat{B}^{2}\widehat{t}_{n}$, where $\widehat{B}^{2}\widehat{t}_{n}^{\,2}$ is bounded above by the worst-case MSE of any linear estimator of the form $\sum_{i=1}^{n}\gamma_{i}Y_{i}$. In particular, there exists a kernel regression estimator that has worst-case MSE (supremum taken over our function class $\mathcal{M}_{\widehat{B}}$) of the order of $\mathcal{O}_{p}(n^{-6/7})$. To make it precise, consider the kernel regression estimator
@@
where $\widehat{B}$ and $\widehat{\sigma}^{2}$ are fit externally, independent of the data $\{X_{i},Y_{i}\}_{i=1}^{n}$. The worst-case bias is given by $\widehat{b}=\widehat{B}\widehat{t}_{n}$, and thus the goal here is to show that $\widehat{t}_{n}^{\,2}/\widehat{V}_{n}=\mathcal{O}_{p}(1)$. Recall the population-level problem with externally fit nuisance parameters:
@@
It suffices to prove that $t_{n}^{*}/\sqrt{V_{n}^{*}}=\mathcal{O}_{p}(1)$, where $V_{n}^{*}=\widehat{\sigma}^{2}\mathbb{E}[\gamma_{n}^{*,2}(X)]/n$, since we have consistency (i.e., that $\widehat{t}_{n}/t^{*}_{n}\rightarrow_{p}1$ and $\widehat{V}_{n}/V_{n}^{*}\rightarrow_{p}1$). It follows from Lemma 12 that $\widehat{B}^{2}t_{n}^{\,*,\,2}+V_{n}^{*}\leq_{p}a_{+}n^{-6/7}$ for some $a_{+}\in(0,\infty)$, for all sufficiently large $n$. On the other hand, it follows from the minimax lower bound (see, e.g.Cheng et al. (1997)) that $\widehat{B}^{2}t_{n}^{\,*,\,2}+V_{n}^{*}\geq_{p}a_{-}n^{-6/7}$ for some $0<a_{-}\leq a_{+}$ for all sufficiently large $n$. Therefore, to claim that $t_{n}^{*}/\sqrt{V_{n}^{*}}=\mathcal{O}_{p}(1)$, it suffices to show that there exists some $c>0$ such that $V_{n}^{*}\geq_{p}cn^{-6/7}$ for all sufficiently large $n$.
@@
Choose $k$ large enough so that $a_{+}k^{-6/7}\leq a_{-}/2$, and then choose $c>0$ small enough such that $ck^{1/7}<a_{-}/2$. This, along with the above display, implies that $R_{m}(\gamma_{m}^{*})\leq R_{m}(\gamma_{n}^{*})\leq_{p}a_{-}m^{-6/7}$, which contradicts the minimax lower bound (we choose $n$ is large enough so that $m=n/k$ satisfies the minimax lower bound $\widehat{B}^{2}t_{m}^{\,*,\,2}+V_{m}^{*}\geq_{p}a_{-}m^{-6/7}$). This contradiction completes the proof. $\square$